# Unexplainability of Artificial Intelligence Judgments in Kant’s Perspective

### v2

@@ Introduction
Because it does not contribute to the content of the judgment, it allows us to explain it in the “Problematical” form even if we do not know what judgment the machine makes. Therefore, even before the advent of explainable AI, it was possible to obtain judgments in the form of “Problematical” only through the SoftMax function. What we want to know through explainable AI may be the characteristics of the entire judgment shown in Table 1, but another way is to artificially consider only the content of the judgment and intentionally fix the modality of the judgment. In other words, it is not that we will see what kind of judgment was made, but that we will treat it as a possibility judgment no matter what judgment was made. People might claim that we can transform some judgments into possibility judgments. For example, if something is said to be inevitably and logically true, the probability is 100%. Can “actually exist,” “necessary,” and “possible” be the same? Treating all judgments as possibility judgments seems problematic.
Limitation of vision-based explainable AI
We may encounter a range of things that cannot be explained when dealing with concepts that have no visual intuitions but whose functions are revealed in vision. I address this topic by beginning with an experiment conducted by (Seo and Lee 2021). To predict the speed of a ball, it is crucial to observe its movement. However, if the camera that is filming it also begins to move in the same manner the ball does, the ball will always appear stationary in the camera’s images. Hence, the position of the camera must also be taken into account. Coordinate information, such as those provided by the global positioning system (El-Rabbany 2002) can be utilized in this regard. Moreover, the camera’s position can be estimated by using fixed background information. If the background screen moves, it means that the camera moves. If it moves only to a minimal degree, the camera moves only minimally, Figure 2 provides an image of this scenario.
Figure 2: A ball, camera, and colorful boxes fixed in the background.
Performance decreased when boxes were eliminated (Seo and Lee 2021). This indicates that the AI used boxes to predict the speed. From this, I conducted additional experiments. Specifically, I seek to demonstrate that the AI utilizes the boxes for information from boxes by employing Grad-CAM (Selvaraju et al. 2017). Grad-CAM focuses on a classification problem, so it cannot be used for the task that predicts the speed itself. Therefore, it is trained to classify class 1 for high speeds based on the average of final speeds in the training set and class 2 for low speeds. When AI predicts speeds higher than the average, I examine where it focuses on its attention.
I employ a combination of a convolutional neural network (CNN) (Lawrence et al. 1997), which processes two-dimensional spatial information (e.g., photos), and a long short-term memory (LSTM) (Hochreiter and Schmidhuber 1997) which processes data obtained along the time axis. This is called a long-term recurrent convolutional network (LRCN) (Donahue et al. 2015). The training data comprise 200 videos, each containing 20 images captured at 0.2-second intervals. The trained model consists of 4 CNN layers, 2 LSTMs, and 1 fully connected layer (LeCun, Bengio, and Hinton 2015). If the values in parentheses represent the input channel, the output channel, the square kernel size, and the stride in order, the parameters for the CNN layers are (3,15,5,2), (15,13,5,2), (13,7,5,2), and (7,3,5,2). The LSTM has two layers and outputs 40 nodes. Finally, the LRCN outputs the speed in the last fully connected layer. The learning rate is set to 1e-4, and Adam (Kingma and Ba 2014) is used. Cross entropy loss function (Wang et al. 2020a) is employed, and the learning epoch is set to 2000. Figure 3 presents the result. I use the activation map of the CNN layer that is first used in the forwarding process, which processes input values.
Figure 3: Where does AI focus?
Figure 3 displays the areas where AI focuses on. The area marked in white is the area that AI focuses on. These white areas are near the balls and the boxes. This indicates that the speed of the ball is predicted in relation to the information from the balls and boxes, as discovered by (Seo and Lee 2021). It is intriguing that even though I only provided the correct classes for the speed of the ball and the video, the AI began to give its attention to the balls and boxes, indirectly utilizing the boxes for self-location.
We could imagine the machine thinking as follows: While tracking the position of the ball and the box, considering that my location has moved by this much, the speed is likely to be at this level. Here, we ask the crucial question: Does the fact that a machine seeks self-location mean that it understands the concept of self? Humans address the problem using self-consciousness. Do machines also have such concepts, based on functionalism? Problems arise from the following. The concepts of dogs and cats have physical intuitions. However, it seems that the concept of self has no such accompanying intuitions. Let us investigate the concepts relating to humans, animals, and AI.
We can think that the structure of human cognition is the same as each other, assuming no illness is present. The cognitive structure itself is completely given and becomes fully experienced. Research seeks to elucidate how cognition functions, leading to investigations of epistemology and logic. In epistemology, Kant argues for the division of self-consciousness into transcendental and empirical aspects. This perspective and other questions persist in philosophy and in the empirical sciences. First of all, several attempts have been made to scientifically examine our self-consciousness. Bodily self-consciousness has been explored (Blanke 2012; Blanke, Slater, and Serino 2015; Ionta et al. 2011; Lenggenhager, Mouthon, and Blanke 2009; Lenggenhager et al. 2007; Serino et al. 2013). In particular, (Blanke 2012) includes attempts to determine the neural mechanisms that underlie self-identification, self-location, and the first-person perspective. The question of whether individuals with autism possess the concept of self and how it can be enhanced has been investigated (Duff and Flattery Jr 2014; Huang et al. 2017; Lind and Bowler 2009; Lombardo et al. 2007; Mitchell and O’Keefe 2008). On the other hand, Kant discussed self-consciousness in the following paragraph (Kant 1855, pp. 81–83).
The I think must accompany all my representations, for otherwise something would be represented in me which could not be thought; in other words, the representation would either be impossible, or at least be, in relation to me, nothing. […] For the manifold representations which are given in an intuition would not all of them be my representations, if they did not all belong to one self-consciousness, that is, as my representations (even although I am not conscious of them as such), they must conform to the condition under which alone they can exist together in a common self-consciousness, because otherwise they would not all without exception belong to me. […] For example, this universal identity of the apperception of the manifold given in intuition, contains a synthesis of representations, and is possible only by means of the consciousness of this synthesis. For the empirical consciousness which accompanies different representations is in itself fragmentary and disunited, and without relation to the identity of the subject. This relation, then does not exist because I accompany every representation with consciousness, but because I join one representation to another, and am conscious of the synthesis of them. Consequently, only because I can connect a variety of given representations in one consciousness, is it possible that I can represent to myself the identity of consciousness in these representations; in other words, the analytical unity of apperception is possible only under the presupposition of a synthetical unity. The thought, “These representations given in intuition, belong all of them to me,” is accordingly just the same as, “I unite them in one self-consciousness, or can at least so unite them;” and although this thought is not itself the consciousness of the synthesis of representations, it presupposes the possibility of it; that is to say, for the reason alone, that I can comprehend the variety of my representations in one consciousness, do I call them my representations, for otherwise I must have as many-coloured and various a self as are the representations of which I am conscious.
When a human solves the task given to the AI, it is necessary that all of the intuitions of all images coming in at every moment belong to a single self-consciousness. Because, if I borrow a phrase from Kant, self-consciousness is conscious of the synthesis of representations. Therefore, self-consciousness is needed to solve the task. And the concept of self is needed for considering self-location.
Animals cannot directly share their experiences with us, so we are in need of evidence to clarify that they are aware of the concepts of self and self-consciousness. Kant also grappled with this difficulty (McLear 2011). Animals seem to have a concept of self. The empirical investigations that have researched this concept in animals have primarily focused on self-location. For instance, place cells, which are found in the hippocampus of mice are known as cells for self-location (Barry and Burgess 2014; Chen et al. 2019; Danjo, Toyoizumi, and Fujisawa 2018; Knierim 2015; Nadel and Eichenbaum 1999). Place cells are activated with respect to the mouse’s location. Specific neurons are activated in specific spatial positions, and these change as the mouse moves. Also, Mice acquire self-location through background information (Chen et al. 2019). The 2014 Nobel Prize in Physiology or Medicine was awarded for the discoveries of the cell. I want to emphasize that animals’ ability to understand self-location is not a mere claim but a well-established fact. These studies might be pieces of evidence that animals process temporal and spatial information, implying the existence of the concepts of self and self-consciousness.
@@
What sets AI apart from humans and animals is that we can create its cognition structure to be similar to our own, as it is a technology we create, not a natural science that must be verified empirically. This leads us to interpret that AI has a cognitive structure that resembles ours, even if only partially, provided that sufficient conditions are provided. These characteristics can lead to attaching the word “Artificial” to an existing concept. Explainable AI involves a long journey that involves identifying the parts of AI that align with human cognition and striving to make it similar to ours. However, AI researchers are not obligated to always make AI explainable.
Based on this generous attitude, previous investigations (Bringsjord et al. 2015; Chella, Frixione, and Gaglio 2008; Zhang and Zhou 2013) have argued that machines have self-consciousness. However, we cannot have any confidence in the concepts that do not have visual intuitions. These concepts could be explainable in other ways, but they cannot be interpreted in a vision-based explainable way. We can only confidently conclude that there is a relationship between these and other parts. Even if vision-based explainable technology develops, it cannot explain anything about concepts that appear in vision but do not have physical intuitions.
We can emphasize that AI is created to perform specific tasks. In natural language processing, the chat of AI is an important topic. When a certain sentence is inputted, the AI outputs a response (Deng et al. 2019; Zhou et al. 2018). Recent technologies have displayed chatbots that can pass the Turing test. Nevertheless, we do not assert that chatbots have self-consciousness. They are just set up to behave that way. Using human chat data, AI is trained to output appropriate sentences. As another example, a generative adversarial network consists of a discriminator and a generator (Creswell et al. 2018; Goodfellow et al. 2014). Here, the generator produces random fake data, and the discriminator is responsible for distinguishing between the real and the fake data. Making them compete with each other, we train the generator to produce data that are as similar to real data as possible. For another example, reinforcement learning studies how an agent adapts to the environment through changing actions and states after dividing a system into an agent and environment (Mazyavkina et al. 2021; Sutton and Barto 2018). Both cases involve humans dividing two groups in advance and making them interact. I doubt whether these computer programs acquire the concept of the self in this process. Even without this concept, they can address problems because we specifically ask them to do this. There is no inherent reason why they must possess these concepts.
Vision, natural language, and definition in Kant’s perspective
When explaining AI judgment, in addition to highlighting pixels, it could be helpful to have a textual output. But we could have doubts. To make a judgment, AI must understand the concepts used in a judgment. Does the machine truly understand the concepts expressed in natural language? In other words, does AI accurately understand the definitions of all the concepts in sentences? Kant said the definition as follows (Kant 1855, pp. 443–444).
A definition is, as the term itself indicates, the representation, upon primary grounds, of the complete conception of a thing within its own limits. Accordingly, an empirical conception cannot be defined, it can only be explained. For, as there are in such a conception only a certain number of marks or signs, which denote a certain class of sensuous objects, we can never be sure that we do not cogitate under the word which indicates the same object, at one time a greater, at another a smaller number of signs. Thus, one person may cogitate in his conception of gold, in addition to its properties of weight, colour, malleability, that of resisting rust, while another person may be ignorant of this quality. We employ certain signs only so long as we require them for the sake of distinction; new observations abstract some and add new ones, so that an empirical conception never remains within permanent limits. […] In the second place, no a priori conception, such as those of substance, cause, right, fitness, and so on, can be defined. For I can never be sure, that the clear representation of a given conception (which is given in a confused state) has been fully developed, until I know that the representation is adequate with its object. But, inasmuch as the conception, as it is presented to the mind, may contain a number of obscure representations, which we do not observe in our analysis, although we employ them in our application of the conception, I can never be sure that my analysis is complete, while examples may make this probable, although they can never demonstrate the fact. […] Consequently, the science of mathematics alone possesses definitions. For the object here thought is presented a priori in intuition; and thus it can never contain more or less than the conception, because the conception of the object has been given by the definition—and primarily, that is, without deriving the definition from any other source.
Every year, hundreds of papers are published in numerous journals to report new findings on neural networks in biology. Although we do not have a complete understanding of neural networks, we have made the definition of neural networks from our limited knowledge. Can such a definition be complete? Nevertheless, most people admit that ANNs have been created that resemble neural networks. In Kant’s context, although it may seem confusing, there exists a sufficiently agreed definition. We can infer that the definition of an empirical concept relies on experience and appears to be closely intertwined with consensus and language utilization. The idea that meaning is determined by the way language is utilized is well expressed in Wittgenstein’s philosophy of language (Travis 1989) and plays an important role in the philosophy of language as well. It is worth noting that natural language processing is also designed with the view that the meaning of a word is determined by the existence of surrounding words (Devlin et al. 2018; Le and Mikolov 2014; Mikolov et al. 2013; Vaswani et al. 2017).
Even if a machine cannot strictly know concepts precisely, if it is true that AI vaguely understands the concepts, then AI understanding them should be admitted. However, it is hard to determine whether AI understands the concepts in the sentences made by AI to the degree that would be generally accepted by humans. For example, it does not seem easy to determine whether AI understands the definitions and meanings of concepts such as “Is,” “Have,” and “This” to a level that humans can accept. In situations where it is questionable whether AI understands each concept, we cannot trust AI judgments shown in sentence form.
Conclusion
If AI is intended to imitate human judgment, it should exhibit the categories of human judgment. According to Kant, four such categories exist. I analyzed AI judgments within each. In the result, it is shown that AI judgments are ambiguous. Because the characteristics of judgments overlap, we can call this AI’s uncertainty. Furthermore, this paper demonstrates that concepts lacking physical intuitions are challenging to elucidate when they are visually represented through their functions. Lastly, I show that even when utilizing natural language processing technology for expression through subject and predicate, which are fundamental components of judgment, it is hard to determine whether AI genuinely comprehends the concepts it employs to a degree acceptable to humans. Therefore, we cannot trust AI judgments.
We have striven to understand AI judgments within the confines of our human intellect. Can AI judgments possess additional characteristics beyond those of human judgments? Even if such judgments can exist, it appears that we lack the means to fully comprehend them. All problems stem from our acknowledgment that AI makes judgments.

### v4

@@ Introduction
Because it does not contribute to the content of the judgment, it allows us to explain it in the “Problematical” form even if we do not know what judgment the machine makes. Therefore, even before the advent of explainable AI, it was possible to obtain judgments in the form of “Problematical” only through the SoftMax function. What we want to know through explainable AI may be the characteristics of the entire judgment shown in Table 1, but another way is to artificially consider only the content of the judgment and intentionally fix the modality of the judgment. In other words, it is not that we will see what kind of judgment was made, but that we will treat it as a possibility judgment no matter what judgment was made. SoftMax acts as an epistemological filter, compelling the interpretation of AI judgments in terms of possibility. As shown in Table 4, when deriving an output from a single event, both the thought content and the modality of judgment can vary. If a further softmax is applied to the resulting judgment, it can transform even an assertoric statement, such as “The object is currently falling,” into a probabilistic form, e.g., “The object might currently be falling with 70% probability.” Thus, even firm judgments can be reframed within a probability distribution. People might claim that we can transform some judgments into possibility judgments. For example, if something is said to be inevitably and logically true, the probability is 100%. However, can “actually exist,” “necessary,” and “possible” be the same? Treating all judgments as possibility judgments seems problematic.
Thus, we observe that unexplainability can exist in any of the four categories of judgment. This paper presents only a few examples, but unexplainability arises in infinitely many cases. Based on the examples provided here, it should be possible to generate many more new examples. Inspired by the uncertainty principle in quantum mechanics (Busch, Heinonen, and Lahti 2007), we can call this AI’s uncertainty. In this case, all forms of judgment overlap. AI judgments have advanced one step beyond mere indeterminacy — where they were simply unknown — by revealing that different forms of judgment are layered together.
In the field of explainable AI research, various technologies are being studied (Adadi and Berrada 2018; Selvaraju et al. 2017). I believe that by developing technologies that provide insights into how AI judgments are made, we can achieve some level of explainability. Computer scientists have begun revealing which part of an image AI focuses on when making judgments. These tools assist us in inferring the subject and predicate. However, neither a subject nor a predicate is explicitly stated. For example, one representative technique involves highlighting which parts of an image led to the judgment that it is a dog. However, there are also areas such as image captioning, where a model generates captions based on a photo, and models that output both an answer and an explanation simultaneously (Sammani, Mukherjee, and Deligiannis 2022). These technologies explicitly generate natural language, producing both subjects and predicates. In doing so, AI begins to expand its mimicry beyond mere judgment (the content of thought) into the realm of reasoning (the act of thinking). The importance of such technology is particularly evident in domains like medical AI, where it is crucial to ensure that the rationale behind a model’s judgment aligns with the reasoning used by human physicians.
Unexplainability related to complete functional implementation
Johannes Hessen, in his textbook on philosophy (Hessen 1947), sharply distinguishes three essential components of thought. Thought consists of the act of thinking, the content of thought, and the object of thought. When we mentally consider something, this corresponds to the act of thinking. For example, in response to the proposition “An elephant has four legs,” one person may not be familiar with elephants and thus unable to visualize one internally, but can still picture an animal with four legs. Another person may mentally imagine an elephant and visualize its four legs. Yet another might picture a running elephant while reconstructing the proposition. In this way, even for a single proposition, the act of thinking occurs differently across individuals. However, the content of thought remains the same. What the proposition expresses is identically conveyed to multiple subjects; this is what we call the content of thought. Finally, the elephant and its legs—as they exist—appear in thought as the object of thought. Psychology addresses the act of thinking; logic is concerned with the content of thought; and ontology, along with the individual sciences, deals with the object of thought. The question of whether the content of thought corresponds to the object of thought always arises—and this is the domain of epistemology. For example, the question “What is truth?” belongs to this field. This summary encapsulates the information provided by Hessen and forms the conceptual foundation for the discussion that follows.
Through NLP technologies, we are now able to simultaneously address the act of thinking, the content of thought, and the object of thought. A model developed in recent research, powered by vast computational resources, has been trained on virtually unlimited amounts of human-generated text, thereby capturing how words are actually used. As a result, these systems now exhibit the ability to functionally replicate aspects of human natural language behavior with remarkable fluency. Such technologies are referred to as large language models (LLMs), with ChatGPT (Achiam et al. 2023) being one of the most well-known examples. In the case of humans, it is impossible to conceive of thought content without the act of thinking—how can one be said to think without engaging in the process of thinking itself? Machines, however, are not bound to follow the human mode of thought. Nevertheless, because they learn from natural language, they indirectly acquire traces of human cognitive acts. For example, they can indirectly learn the act of thinking by learning how the flow of human thought is expressed through language. Natural language examples showing whether the content of thought matches its object allow for indirect learning about the object of thought. In keeping with the earlier perspective, we set aside ontological questions and focus solely on the functional aspects. That is, LLMs do not truly understand or possess the meanings of such words; instead, they learn how these words are used in relation to others, and are thus able to partially reproduce their functional behavior.
Even in the imitation of thought through NLP, we inevitably confront the unexplainability of AI, which arises from the intrinsic nature of language. According to Kant, a definition is always incomplete and can only express a concept to the extent that it is distinguishable (Kant 2024, pp. 212–213). In short, one person may define “gold” by its color, weight, and chemical properties, while for another, the definition may omit weight and chemical characteristics. Since it is impossible to construct a fixed and rigorous definition with complete precision, the meanings of words cannot be permanently fixed. We do not possess a complete definition, and without it, we cannot fully understand the complete function of a given word. There is no criterion by which the complete functional implementation of a word can be assessed. In other words, there exists an unexplainability regarding the function of the word.
Furthermore, although the function remains only partially implemented, there are cases where it seems as though the function of a word were fully realized. Through exposure to vast amounts of text, LLMs can freely use the concept of “I.” In this context, AI does not possess the same identity as a human being in terms of physical, chemical, or any other existential aspects; rather, it appears to have acquired only the ability to manipulate the concept of “I” to the extent that is functionally necessary in natural language. But can we detect whether the concept of the self has been fully implemented functionally through natural language? Even when we pose questions through natural language, we can observe that AI responds quite well, having been trained on vast amounts of data. However, this is still far from a complete functional implementation. I will demonstrate that there also exists a vision-based component of the self through a simple experiment used by (Seo and Lee 2021).
Figure 1: A ball, a camera, and colorful boxes are fixed in the background. The task involves both the camera and the ball moving, and the goal is to match the ball’s velocity. The white areas indicate the regions where the AI is focusing its attention.
I employ a long-term recurrent convolutional network (LRCN) (Donahue et al. 2015) and Grad-CAM (Selvaraju et al. 2017). LRCN is an algorithm that learns from video input, where two-dimensional images are acquired sequentially over time and stacked. Grad-CAM is an algorithm that provides visual evidence for why an AI system made a particular decision. The target values were provided as a classification problem, aiming to predict whether the ball was faster or slower than the average speed. Details related to the training process are omitted here, as they are not essential to the current discussion.
@@
Of course, it is premature to conclude that AI has a function of the self. It may be more appropriate to acknowledge that some degree of self-function is present when, through extensive experience, one comes to understand that it is not tied to any particular color or shape. In natural language, we observe that LLMs can seemingly use the concept of “I” with complete fluency. LLMs can even provide a fluent explanation of the vision-based self-functions using sentences. However, LLMs still lack vision-based self-functions. This demonstrates that the adequate realization of a particular function may obscure the incompleteness of other functions.
Caution: functional similarity is not ontological identity
To be clear, this does not mean that the self in animals and machines is identical in existence to the human self, but rather that certain functions of the self are similar.